# DOSA-MO: Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems improves biomarker discovery

### 27 Dec 2023 v1

@@ 3 Results
We compare 6 different instantiations of the DOSA-MO by varying the regression model used for the adjustment, including the zero regression model, equivalent to not applying any adjustment. The regression models receive as input 3 high level features of the solutions that are correlated with the overestimation: the original fitness, its standard deviation, and the cardinality of the feature set (number of genes). For an example of correlations between the high level features and the overestimation see Supplementary Section 2. We use NSGA3-CHS as wrapped model. The tests are repeated with 8 different combinations of validation type, datasets, objectives, and classification inner model. For each combination we report two measures for the accuracy of the fitness estimation: the MO performance error (Section 3.1.1) and the Pareto delta (Section 3.1.2). Additionally, we report a measure of overall performance of the optimizers: the CHV (Section 3.2). In Section 3.3 we show in detail the performance of the best solution sets from the external validation.
@@ 3.1 Overestimation in feature selection for biomarker discovery
We measured the error of the performance estimates with two methods: the MO performance error and the Pareto delta, both of which are detailed in the methods section. In summary, the MO performance error quantifies the discrepancy between the HV computed on the train performance and the CHV. On the other hand, the Pareto delta is a metric that captures differences between expected performance and actual performance on new samples, focusing on the more granular level of individual solutions.
@@ 3.1.1 Comparative analysis using multi-objective performance error
Figure 1: MO performance error results for internal-external CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification, and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent standard deviation between folds.
The MO performance error evaluates the DOSA-MO algorithm’s capability in reducing the performance estimation error. The MO performance error quantifies the discrepancy between two key metrics that each summarize a solution set’s fitness in a single value: the HV computed from training performance and the CHV (see Section 2.1.1). We report this error across various experimental setups. The MO performance error outcomes from various regression-based adjusters are reported in Fig. 1.
Significantly, ptree and RFReg often outperform other overestimation predictors in terms of MO performance error. Therefore, for users prioritizing the reduction of MO performance error, ptree and RFReg regressors emerge as particularly effective choices. In comparison, SVR-based estimators of overfitting appear to be the least effective when considering MO performance error. Moreover, it is interesting to note that ptree and RFReg regressors perform particularly well in multi-cohort biomarker discovery and validation scenarios. In these scenarios, an external dataset (SCAN-B) is used to evaluate the models (biomarker sets), with their performance adjusted based on what has been learned from the initial TCGA-BRCA cohort.
@@ 3.1.2 Comparative analysis using the Pareto delta
Figure 2: Pareto delta results for internal-external CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent standard deviation between folds.
The Pareto delta measure (Section 2.1.2) considers the difference between a model’s predicted and actual test performance for each individual solution. This makes it particularly relevant and more informative than the MO performance error in the common situation when users are presented with multiple solutions but will ultimately select only one. A comparison of the Pareto delta measure across various experimental setups is detailed in Fig. 2.
In the DOSA-MO framework, ptree and RFReg regressors consistently outshine others when reducing performance estimation error, aggregated on a solution by solution basis with the Pareto delta. Interestingly, while no significant differences are noted in MO performance error between the dummy model and the unadjusted optimizer, the Pareto delta metric reveals that all kidney cancer setups with proper regressors have a lower Pareto delta compared to the unadjusted optimizer. rSVR performs better than the zero regressor in all scenarios except one. RFReg and ptree always yield lower Pareto deltas than the zero regressor, including in breast cancer setups, highlighting their promise in both overestimation measures.
@@ 3.2 Evaluating DOSA-MO’s effectiveness in MO feature selection
Figure 3: CHV results for internal-external CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent standard deviation between folds.
The previous results highlight DOSA-MO’s success in reducing performance estimation errors in ML-driven feature selection for biomarker discovery. This section evaluates if DOSA-MO’s dual-stage framework also enhances the quality of the produced feature set, that is the final output for biomarker selection models considered for clinical validation. To assess DOSA-MO’s impact on the overall feature selection process, we calculated the CHV for each experimental setup (see Fig. 3).
Tree-based regressors effectively reduce estimation errors but do not significantly improve the quality of solution sets generated by the MO optimizer. This indicates that using these regressors for overestimation prediction may not always enhance solution sets, despite reducing quality estimation errors. Interestingly, dummy and SVR-based MO optimizers consistently outperform the CHV of unadjusted models, highlighting the effectiveness of overestimation prediction in DOSA-MO for superior feature selection outcomes. Specifically, SVR outperforms the zero model in all kidney setups and most breast setups, except in scenarios involving the overall survival objective. Notably, SVR yields optimal results in setups excluding survival prediction, while the dummy model excels in setups that include survival prediction.
3.3 Advancing biomarker discovery using multi-cohort studies.
In this section, we examine a case study that utilizes two distinct cohorts, TCGA-BRCA and SCAN-B, for external validation in biomarker discovery. The approach focuses on learning models and overestimation during biomarker identification within one cohort (like TCGA) and then applying this knowledge to another cohort (such as SCAN-B). This shows the reliability of biomarker discovery across different cohorts.
(a) (b) (c)
As previously shown, in the external validation for classification of breast cancer subtypes the DOSA-MO with SVR as fitness adjustment regression model and SVM as inner classifier provides the best overall performance measured with the CHV, while, keeping the SVM as inner classifier, using the RFReg it is possible to obtain the more accurate prediction of the performance on a new cohort, with respect to both MO performance error and Pareto delta. The effect of the adjustment can be appreciated by looking at the solution sets presented in Fig. 4. The balanced accuracy expected by the optimizer (”internal cv” in the figures) is monotonically increasing with the number of features because the solutions returned by the algorithm are non-dominated. The estimation error, i.e. the vertical distance between the balanced accuracy expected by the optimizer using only the training samples and the balanced accuracy measured on the external dataset, is lower when using the DOSA-MO.
@@ 4 Discussion
We presented a new wrapper MO optimizer capable of adjusting the performance measures for the overestimation that is due to multiple model validations, improving the solution set that is produced. It has been validated on multiple tanscriptomics configurations related to kidney and breast cancer, including external validation.
Since to the best of our knowledge there was no measure of error in the performance estimation for MO solution sets, we proposed two such measures. According to both of them, the DOSA-MO provided improved performance estimates in all the considered scenarios when a regression model based on decision trees was used to predict the overestimation.
Even the dummy regression model that learns just the weighted average of the overstimation, when used for adjusting the fitness, allows to obtain a better overall performance than the unadjusted optimizer in 7 out of 8 experimental setups, as measured by the CHV. The SVR regression model also provides for a better CHV than the unadjusted optimizer in 7 out of 8 setups. The dummy regression model is the best in the 3 setups that include survival prediction, while the SVR is the best in the 5 setups that have subtype classification and economy of features as objectives.
The rSVR regression model selects the hyperparameters for the regularization strength of the SVR using random search. It performed poorly in our tests. A regularization that works well while cross-validating on solutions obtained by the same unadjusted optimizer, is not as effective when the GA runs with a bigger population, more generations, and the adjustment applied to the evaluation of the individuals. More conservative regression models, like the dummy, or the SVR with its strong default regularization, have proven more effective in our tests.
Predicting the overestimation is more complex when the adjustment to the fitness is applied during the optimization as in our approach. The regression models are trained on sample distributions, but then they apply the adjustment results to different solutions affecting their selection. We cannot expect for the regression models to be as precise in areas of the solution space that have not been explored when creating the training set. This applicability domain issue brings an added difficulty.
Adjusting the fitness impacts the exploration of the optimizer, and the explored solutions will have a different distribution from the one on which the regressors were trained, even in the areas included in the applicability domain. This is a moving target problem, often solved in other cases by optimizing in small steps, for example when applying back-propagation to artificial neural networks across multiple epochs. It is not practical to run many iterations of computationally expensive MO GAs, but strategies involving faster algorithms may be explored. Since the solutions can potentially still be used for training the regression models also in the following steps, techniques may be mutuated from on-line learning algorithms. More research is needed to improve how the applicability domain and moving target problems are addressed. The proposed algorithm can be applied to any data and problem and has been experimentally validated on cancer transcriptomics data for subtype classification and survival prediction. Further experimental validation is needed to assess its performance in other domains.

### 29 Feb 2024 v3

@@ 3 Results
(a) (b) (c)
We compare 6 different instantiations of the DOSA-MO by varying the regression model used for the adjustment, including the zero regression model, equivalent to not applying any adjustment. The regression models receive as input 3 meta-features of the solutions correlated with the overestimation: the original fitness, its SD, and the feature set size (number of genes). For an example of correlations between meta-features and overestimation see Supplementary Section 2.1. We use NSGA3-CHS as wrapped model. The tests are repeated with 8 different combinations of validation type, datasets, objectives, and classification inner model. For each combination we report two measures for the accuracy of the fitness estimation: the MOPE (Section 3.1.1) and the $P_{\Delta}$ (Section 3.1.2). Additionally, we report a measure of overall performance of the optimizers: the CHV (Section 3.2). A focus on the best solution sets from the external validation is shown in Fig. 2 and discussed in Supplementary Section 2.2.
@@ 3.1 Overestimation in feature selection for biomarker discovery
We measured the error of the performance estimates with two methods: the MOPE and the $P_{\Delta}$. The MOPE quantifies the discrepancy between the HV computed on the train performance and the CHV. In contrast, the $P_{\Delta}$ captures differences between expected performance and actual performance on new samples, focusing on the more granular level of individual solutions.
@@ 3.1.1 Comparative analysis using multi-objective performance error
Figure 3: MOPE results for internal k-fold CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification, and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent SD between folds.
The MOPE evaluates the DOSA-MO algorithm’s capability in reducing the performance estimation error (Section 2.4.1). The MOPE outcomes from various regression-based adjusters are reported in Fig. 3.
Significantly, ptree and RFReg often outperform other overestimation predictors. Therefore, for users prioritizing the reduction of MOPE, ptree and RFReg regressors emerge as particularly effective choices. SVR-based estimators of overfitting appear to be the least effective. ptree and RFReg regressors perform particularly well in external validation scenarios, where an external dataset (SCAN-B) is used to evaluate the models (biomarker sets), that were selected according to their adjusted performance learned from the initial TCGA breast cohort.
@@ 3.1.2 Comparative analysis using the Pareto delta
Figure 4: $P_{\Delta}$ results for internal k-fold CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent SD between folds.
The $P_{\Delta}$ measure (Section 2.4.2) considers the difference between a model’s predicted and actual test performance for each individual solution. This makes it particularly relevant and more informative than the MOPE in the common situation where users are presented with multiple solutions but will ultimately select only one. The $P_{\Delta}$ across the experimental setups is shown in Fig. 4.
ptree and RFReg regressors consistently outshine others when reducing the $P_{\Delta}$. Interestingly, while no significant differences are noted in MOPE between the dummy model and the unadjusted optimizer, all kidney cancer setups with proper regressors have a lower $P_{\Delta}$ compared to the unadjusted optimizer. rSVR performs better than the zero regressor in all scenarios except one. RFReg and ptree always yield lower $P_{\Delta}$s than the zero regressor, including in breast cancer setups, highlighting their promise in both overestimation measures.
@@ 3.2 Evaluating DOSA-MO’s effectiveness in MO feature selection
Figure 5: CHV results for internal k-fold CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent SD between folds.
The previous results highlight DOSA-MO’s success in reducing performance estimation errors in ML-driven feature selection for biomarker discovery. It is also important to evaluate if DOSA-MO enhances the quality of the produced feature sets: the final output for biomarker selection models considered for clinical validation. To assess the impact on the overall feature selection process, we calculated the CHV for each experimental setup (see Fig. 5).
Tree-based regressors reduce estimation errors but do not significantly improve the quality of solution sets generated by DOSA-MO. This indicates that using these regressors for overestimation prediction may not always enhance solution sets, despite reducing quality estimation errors. Interestingly, dummy and SVR-based MO optimizers consistently outperform the CHV of unadjusted models, highlighting the effectiveness of overestimation prediction for superior feature selection outcomes. Specifically, SVR yelds the best CHV in all setups except in scenarios involving the overall survival objective, while the dummy model excels in setups that include survival prediction.
@@ 4 Discussion
We introduced DOSA-MO, a novel wrapper MO optimizer designed to adjust performance measures for overestimation in ML-driven MO problems, improving the solution sets that are produced. This approach has been successfully validated across multiple transcriptomic-based datasets commonly used for biomarker discovery in cancer subtype classification and survival prediction. Our benchmarking extends to a scenario where biomarker models, trained in one population-based cohort (such as TCGA), are applied to classify cancer subtype in a second cohort (such as SCAN-B). This method effectively enables biomarker discovery and validation across multiple cohorts. Additionally, we have developed two innovative measures for evaluating the performance of MO algorithms: the MOPE and the $P_{\Delta}$. According to both metrics, the DOSA-MO algorithm demonstrated improved performance estimates in all tested cases, particularly when using decision tree-based regression models for predicting overestimation.
Our study found that even a basic regression model, which learns solely the weighted average of overestimation and is used for fitness adjustment, resulted in improved overall performance compared to the unadjusted optimizer in 7 out of 8 experimental setups, as indicated by the CHV metric. Likewise, the MO optimizer directed by the SVR model, which predicts overestimations, surpassed the unadjusted optimizer in 7 out of 8 setups, affirming its effectiveness. Notably, the MO optimizer guided by the simpler dummy regression model excelled in the three setups involving survival prediction, while the SVR model proved superior in the five setups focused on subtype classification and feature set size. The rSVR regression model, that selects the hyperparameters for the regularization strength of the SVR with random search, utilized for predicting overestimation, underperformed with respect to dummy and SVR regression models. A regularization that works well while cross-validating on solutions obtained by an unadjusted optimizer, is not as effective when the GA runs with a bigger population, more generations, and the adjustment is applied to the evaluation of the individuals.
Predicting the overestimation is more complex when the adjustment to the fitness is applied during the optimization as in our approach. The fitness adjustments alters the optimizer’s exploration path, leading to a divergence in the solution distribution from the one used to train the regressors for overestimations, even within their applicability domain. This creates a “moving target” problem, commonly addressed in other contexts, like artificial neural networks, through incremental optimization across multiple epochs. Although executing more computationally intensive MO GAs is possible, further research is essential to enhance our understanding and address the challenges of the applicability domain and the moving target problem in these contexts.
DOSA-MO is versatile and can be applied to any ML problem, including the ones requiring a MO evaluation framework, making it particularly well-suited for biomarker discovery in high dimensional molecular datasets. It has been experimentally validated with cancer transcriptomics data, demonstrating its applicability and effectiveness in this domain.
Funding
This work was supported by the Academy of Finland [grant agreements 336275, 332510, and 358037], the Jane and Aatos Erkko Foundation [210026], and the Sigrid Jusélius Foundation.
Acknowledgments
The computational analyses were performed on servers provided by the UEF Bioinformatics Center, University of Eastern Finland, Finland.